# AGN-heated dust revealed in “Little Red Dots”

### v1

@@
Little Red Dots (LRDs) are a puzzling population of extragalactic sources whose origin is highly debated. In this Letter, we stack NIRCam, MIRI and ALMA images of a large and homogeneously-selected sample of LRDs from multiple JWST Legacy fields. We find clear evidence for hot-dust emission, seen as a rising mid-infrared continuum up to rest-frame 3m, that is best explained by a standard dusty AGN structure. This scenario holds regardless of whether the Optical/Ultraviolet (UV) continuum is stellar or AGN-dominated. Either ways, we argue that AGN-heated dust must have a geometry that leaves at least part of the Optical/UV continuum and broad-line region emission unobscured, which explains the observed blue UV slopes and prominent Balmer features. Lack of detections in the deep stack of X-ray images suggests that Compton-thick (31024 cm-2) gas obscuration is common, and likely confined within the dust sublimation radius (sub0.1 pc). While this gas-dust displacement is in line with recent models (e.g. the “Black Hole Star”), how pre-enriched hot dust can form around a nearly-pristine black hole environment remains to be explained.
@@
The James Webb Space Telescope (JWST) has ushered extragalactic astronomy in a transformative era, offering unprecedented access to the high-redshift universe. One of the most puzzling discoveries is a population of sources dubbed as “Little Red Dots” (LRDs, e.g. Matthee et al. 2024, Greene et al. 2024, Labbe et al. 2025). LRDs are extremely compact (often unresolved at 100 pc scales) and show red optical colors combined with blue UV slopes, resulting in a characteristic “V-shaped” spectral energy distribution (SED), with a sharp turnover near the Balmer break at rest Å (e.g. Setton et al. 2024, Hainline et al. 2025). Remarkably, LRDs appear to be two orders of magnitude more abundant than luminous quasars at similar redshifts (Kokorev et al. 2024, Akins et al. 2024), making them accessible even in narrow-field JWST surveys.
However, an observational consensus of the physical nature of LRDs is still lacking. The red optical colors and broad emission lines point toward obscured Active Galactic Nuclei (AGN) activity (Inayoshi et al. 2020, Volonteri et al. 2021, Ji et al. 2025). In this case, the AGN continuum would be moderately reddened by dust (– mag; Casey et al. 2024; Akins et al. 2024; Wang et al. 2025). Yet, this interpretation is apparently at odds with their extreme X-ray and radio weakness implied by the lack of stacked X-ray (Ananna et al. 2024; Yue et al. 2024; Maiolino et al. 2025) and radio emission (Akins et al. 2024; Gloudemans et al. 2025). To alleviate this tension, some theoretical models postulate LRDs as supermassive black holes (SMBHs) accreting at super-Eddington luminosities and enshrouded in a turbulent, dust-free, and highly dense atmosphere: the “BH-star” scenario. This explains the strong Balmer break and the red continuum through pure gas opacity, and the broad Balmer lines via resonant electron scattering (e.g. Naidu et al. 2025; Inayoshi & Maiolino 2025; Rusakov et al. 2025; de Graaff et al. 2025). Similar gas conditions are postulated in other models, with the presence of a metal-free halo surrounding a pre-collapse “super-massive star” (Nandal & Loeb 2025), or a late-stage “Quasi star” (Begelman & Dexter 2025). All these models share the absence of prior metal enrichment and the presence of a gas-dense, dust-free region at sub-pc scales. So far, this dust paucity is supported by the flat LRD SED observed beyond rest 1.6 m (e.g. Williams et al. 2024; Carranza-Escudero et al. 2025), in contrast to the rising mid-infrared (MIR) continuum expected in obscured AGN, conveying the idea that MIR emission in LRDs is dominated by dust-poor stellar populations and not AGN dust (Williams et al. 2024; Pérez-González et al. 2024, Leung et al. 2024). In addition, the non-detection from deep ALMA stacking poses a stringent dust mass limit of Mdust106 M⊙ at 6 (Casey et al. 2025; see also Chen et al. 2025; Setton et al. 2025). Nevertheless, we caution that previous NIRCam/MIRI SED studies of LRDs might have been biased towards bright objects (e.g. Labbe et al. 2024; Barro et al. 2024; Chen et al. 2025; de Graaff et al. 2025; Wang et al. 2025), or possibly affected either by individual MIRI non-detections (e.g. Pérez-González et al. 2024) or by stacking peculiar LRD sub-samples (e.g. -dropouts; Williams et al. 2024), whose SED could intrinsically differ from that of “typical” LRDs. To explore this possibility, in this Letter we stack a large and homogeneously-selected sample of LRDs with broad NIRCam/MIRI coverage. Throughout, we assume a Chabrier (2003) IMF and a flat CDM cosmology with = 70 km/s/Mpc, , and .
2 Photometric selection of LRDs
We select LRDs from the latest compilation of Kocevski et al. (2025), consisting of 341 photometrically-selected LRDs in the redshift range 2–11, and retrieved from several JWST/NIRCam surveys. These LRDs were homogeneously selected to display compact morphologies and “V-shaped” continuum from spectral slope fitting between blueward and redward of the rest-frame 3645 Å (i.e. close to the Balmer break). These criteria enable a self-consistent search for LRD candidates with red optical and blue UV colors over a wide redshift interval (Hainline et al. 2025), although we acknowledge that it does not guarantee an homogeneous physical selection of LRDs.
We replace photometric redshifts () with NIRSpec-based spectroscopic redshifts () taken from the public Dawn JWST Archive (DJA; Brammer 2023b; de Graaff et al. 2024; Heintz et al. 2024), whenever of high-quality (grade=3, see archive). This increases the spectroscopic fraction from 11% (39/341) to 29% (100/341). The mutual agreement between and estimates is generally good, with only five sources showing an offset /(1+)0.2.
Finally, we restrict ourselves to the NIRCam fields with (at least partial) MIRI coverage in two or more bands (including 18m and/or 21m; see Table 1), which yield the strongest IR-SED constraints. Our final sample counts 302 LRDs taken from the following surveys: the Cosmic Evolution Early Release Science Survey (CEERS; Finkelstein et al. 2023) in the EGS; the JWST Advanced Deep Extragalactic Survey (JADES; Eisenstein et al. 2023) in the GOODS-S, complemented by MIRI imaging from the Systematic Mid-infrared Instrument Legacy Extragalactic Survey (SMILES; Rieke et al. 2024; Alberts et al. 2024); the Public Release IMaging for Extragalactic Research (PRIMER111https://primer-jwst.github.io/) survey in the COSMOS and UDS fields. Out of 302 LRDs, 163 (54%) are covered by MIRI 7.7m, 126 (42%) also by MIRI-18m, and 85 (28%) have a robust . Moreover, we exploit archival ALMA/B6 (1.1 mm) data from the combined A3COSMOS+A3GOODSS surveys (Liu et al. 2019; Adscheid et al. 2024; Magnelli et al. 2024), which are available for a total of 70 LRDs in the COSMOS and GOODS-S fields. The number of sources split down by field, photometric coverage and spectroscopic redshifts is summarized in Table 1.
3 Stacking results
We perform median stacking over all available NIRCam bands (7) and all MIRI bands (2 in PRIMER, 7 in CEERS and 8 in SMILES), whenever available, at the optical position of all LRDs. We use a fixed pixel scale of 0.06′′, and we extract fluxes from a circular aperture with fixed radius =0.25′′. ALMA mean stacking, instead, is performed in the -plane, in order to homogeneously combine archival data at different frequency, angular resolution and sensitivity. Further details of the stacking methodology are given in Appendix A. We refer the reader to Appendix B, to verify the robustness of our stacking results against possible biases. In Fig. 1 we show the stacked cutouts (3′′3′′) in each band, sorted by increasing wavelength. For comparison, the dashed white circle marks the corresponding PSF FWHM. As MIRI images are strongly oversampled, to aid visualization we smooth them with a Gaussian kernel, adopting a larger radius for larger PSF FWHM (see caption of Fig. 1). The stacked ALMA/B6 image is obtained in the -plane and imaged at 0.1′′/px scale. Except for MIRI/f2550w and ALMA/B6, all other median stacks lead to a detection with signal-to-noise ratio (S/N) greater than 3. We stress that all detections are consistent with being unresolved, as their extent is smaller or comparable with the PSF FWHM of the corresponding band (white circles in Fig. 1). All fluxes, errors, rms of the stacked images are listed in Table 2 along with the effective number of sources stacked in each band.
4 Discussion and Conclusions
4.1 SED-fitting with vs without AGN
In order to interpret the stacked photometry of LRDs, we perform SED-fitting using the Code Investigating GALaxy Emission (CIGALE; Boquien et al. 2019). We fix the input redshift to the median value of the LRD sample (6.2). In order to account for the broad redshift distribution of the sample (mostly at 4.58; see Fig. 4) and the correspondingly wide rest-frame range sampled by each filter, we convolve the filter response curves at 6.2 with the band-specific redshift distribution of the underlying LRD sample. Then, we run CIGALE both with and without AGN templates to test the relative significance of the AGN component. In both SED runs, the library of stellar templates is taken from Bruzual & Charlot (2003), including a delayed star formation history with random bursts. Energy balance is assumed between the dust-absorbed UV radiation and the reprocessed IR emission, modeled via a dust attenuation curve from Charlot & Fall (2000), and dust templates from Dale et al. (2014). The AGN templates come from the skirtor module (Stalevski et al. 2012; 2016), which includes emission from an accretion disk, a dusty torus and polar dust. A comprehensive list of the SED-fitting parameters is given in Appendix C and D. In Fig. 2, we display all stacked fluxes (or 3 upper limits) overlaid to the total best-fit SED (black line) obtained from the run without AGN (top panel) or with AGN (middle panel).
We caution the reader that our SED-fitting analysis is not primarily aimed at interpreting the rest optical or far-IR properties of LRDs. For instance, CIGALE assumes a single-zone obscuration between optical, UV and IR light, which might be oversimplistic in case of no co-spatiality (see Sect. 4.2). Nonetheless, regardless of the origin of Optical/UV emission in LRDs, we argue that AGN-heated dust is needed to reproduce the monothonic increase up to rest-frame 2–3m, which pure dust-attenuated stellar light fails to match. This is further confirmed by our ALMA upper limit, which is very similar to that derived by Casey et al. (2025) (see also Xiao et al. 2025; Setton et al. 2025), which translates into Mdust106 M⊙ for T100 K. Indeed, the best-fit SED without AGN strongly overshoots the ALMA upper limit when attempting to reproduce the power-law-like SED at rest 1–3m. While this could be resolved by lifting the energy balance implemented in CIGALE, even ignoring the upper limits (at f2550w and ALMA/B6), a simple F-test comparing the two reduced values favours the AGN solution at 95% significance.
4.2 Are LRDs typically dust-obscured AGN?
By considering the best-fit solution with AGN, the rest-frame 6 m luminosity of the AGN torus can be used as a proxy for the intrinsic AGN X-ray (rest [2–10] keV) luminosity (X,intr (e.g. Stern 2015). From 6μm=(3.80.6)1043 erg/s, we derive X,intr=21043 erg/s, which nicely matches that expected from the best-fit AGN bolometric luminosity, if assuming standard X-ray bolometric corrections (e.g. Lusso et al. 2012). By stacking the deepest X-ray images of the 44 LRDs falling inside the Chandra Deep Field South (Luo et al. 2017), we obtain a 3 [2–10] keV luminosity limit X1042 erg/s at =6.2 (Comastri et al. subm). In order to match it with the X,intr expected from the AGN SED, the hydrogen column density has to be NH31024 cm-2, implying Compton-thick gas obscuration (Ananna et al. 2024; Yue et al. 2024; Maiolino et al. 2025).
Given the SED-based estimate of the AGN bolometric luminosity, erg s-1 and the best-fit stellar mass from the AGN solution (M⋆2.5109 M⊙; i.e. 2 smaller than from a pure-galaxy fit), we emphasize that, if LRDs are highly-accreting AGN (Eddington ratio 10%), the BH-to-galaxy mass ratio MBH/M⋆1:100, implying strongly overmassive BHs compared to local scaling relations (see also Pacucci et al. 2023; Juodžbalis et al. 2024; Maiolino et al. 2024). Moreover, our estimate corresponds to a dust sublimation radius of 0.1 pc at K (Barvainis 1987), meaning that any dusty AGN structure should begin at scales pc from the central SMBH. In the bottom panel of Fig. 2, we try to decompose the AGN-dust template with three multi-temperature greybody curves (with free to vary). Our rising MIRI SED up to rest-frame 2–3m indicates hot dust at K, though it could be even hotter if optically-thick. Instead, the bulk dust mass (Mdust3104 M⊙) comes from colder polar AGN dust at 50 K (though not constrained by our data). While a dusty structure could induce X-ray obscuration, it would also strongly suppress the Optical/UV continuum, in tension with the blue UV slope and strong Balmer features that imply relatively modest dust attenuation (; Akins et al. 2024; Casey et al. 2025).
This apparent contradiction can be resolved if X-ray absorption in LRDs occurs within , in a gas-dense and dust-free region. This gas-dust displacement is formally consistent with the recently-proposed models of Quasi-star (Begelman & Dexter 2025), BH–star (Naidu et al. 2025), and super-massive star (Nandal & Loeb 2025) scenarios. However, the presence of pre-enriched dust around a first-generation (or metal-poor) star is hard to justify. Moreover, the normalized BH–star SED (green line in Fig. 2, bottom panel) appears to flatten out at rest-frame 1 m, contrary to the rising MIRI fluxes.
Our findings might fit better with an alternative hypothesis suggested by previous studies (Setton et al. 2024; Maiolino et al. 2025; D’Eugenio et al. 2025; Ji et al. 2025; Inayoshi & Maiolino 2025), in which X-ray obscuration arises from Compton-thick gas clouds within the BLR (at sub, i.e. dust-free), which covers the compact X-ray corona without suppressing the broad Balmer lines. Yet, the origin of the Optical/UV continuum remains open. Although we caution that LRDs might not consist of a single population, we speculate two simple scenarios in Fig. 3, depending on what powers the Optical/UV continuum in LRDs. In either cases, the bulk of Optical/UV continuum is not dust-attenuated, despite the presence of hot AGN dust, and the outer BLR is directly visible along the line of sight. (i) If the Optical/UV comes from stellar light (Fig. 3, left), the LRD resembles a “type-2” AGN configuration, in which the torus covers the AGN accretion disk, which would otherwise outshine the host. In case of high dust clumpiness, mildly-attenuated light from the outer accretion disk could add to the optical emission. (ii) Instead, if the Optical/UV comes from AGN light (Fig. 3, right), the LRD resembles a “type-1” unobscured AGN configuration, though the origin of the sharp Balmer break remains unclear. Fainter Optical/UV light from kpc-scale star formation cannot be ruled out (e.g. Rinaldi et al. 2025).
In summary, we find clear evidence for a rising MIRI continuum up to 3m, which we interpret via SED fitting as hot (8205 K) AGN dust, regardless of whether the Optical/UV continuum is stellar- or AGN-dominated, or a combination thereof. Despite the presence of hot dust, these LRDs selected via V-shaped and compactness criteria must be not strongly dust-obscured. We acknowledge that such homogeneous selection does not necessarily translate into a single LRD population with common physical properties. Our findings mainly provide en empirical benchmark for the presence of AGN-heated dust in the majority of LRDs, or (at least) in a significant sub-population of them. For instance, if LRDs are standard AGN, we speculate that the above selection criteria might be biased against the most dust-obscured (“type-2”) LRD analogs, which could be recovered by lifting the blue-UV slope requirement.
Acknowledgements.
ID acknowledges funding by the European Union – NextGenerationEU, RRF M4C2 1.1, Project 2022JZJBHM: ”AGN-sCAN: zooming-in on the AGN-galaxy connection since the cosmic noon” - CUP C53D23001120006. (Some of) the data producs presented here were retrieved from the DJA, an initiative of the Cosmic Dawn Center (DAWN), which is funded by the Danish National Research Foundation under grant DNRF140.References
- Barvainis (1987) Barvainis, R. 1987, ApJ, 320, 537
- Brammer (2023a) Brammer, G. 2023a
- Brammer (2023b) Brammer, G. 2023b
- Bruzual & Charlot (2003) Bruzual, G. & Charlot, S. 2003, MNRAS, 344, 1000
- Chabrier (2003) Chabrier, G. 2003, PASP, 115, 763
- Charlot & Fall (2000) Charlot, S. & Fall, S. M. 2000, ApJ, 539, 718
- D’Eugenio et al. (2025) D’Eugenio, F., Maiolino, R., Perna, M., et al. 2025, arXiv e-prints, arXiv:2503.11752
- Heintz et al. (2024) Heintz, K. E., Watson, D., Brammer, G., et al. 2024, Science, 384, 890
- Inayoshi & Maiolino (2025) Inayoshi, K. & Maiolino, R. 2025, ApJ, 980, L27
- Juodžbalis et al. (2024) Juodžbalis, I., Maiolino, R., Baker, W. M., et al. 2024, Nature, 636, 594
- Luo et al. (2017) Luo, B., Brandt, W. N., Xue, Y. Q., et al. 2017, ApJS, 228, 2
- Stern (2015) Stern, D. 2015, ApJ, 807, 129
Appendix A Stacked samples and fluxes
Here we provide a more detailed outline of the input images and stacking technique used in this work.
Image and noise maps are retrieved from the corresponding mosaics of each field (see below). In this step, we discard narrow or medium-band filters (whenever available) to mitigate potential flux boosting in the case of strong emission lines.
For the GOODS-S, we retrieve NIRCam and MIRI images from the JADES and SMILES surveys publicly available on the MAST222SMILES: https://archive.stsci.edu/hlsp/smiles; JADES: https://archive.stsci.edu/hlsp/jades archive. While SMILES maps are already at 0.06′′/px, JADES maps (with a native pixel scale of 0.03′′) are re-scaled to 0.06′′/px for consistency with the other fields.
For CEERS and PRIMER-UDS, which are unavailable on MAST, we leverage the DJA (Brammer 2023a; Valentino et al. 2023), where the full reduced mosaics are publicly available in each NIRCam and MIRI band333DJA full reduced mosaics in each band are taken from https://s3.amazonaws.com/grizli-v2/JwstMosaics/v7/index.html, all with a pixel scale of 0.04′′ and in units of [10 nJy]. We do a resampling of all mosaics to 0.06′′/px for consistency with SMILES and PRIMER-COSMOS maps, while preserving astrometry and total flux over a given aperture. For PRIMER-UDS, we also merge the East and West mosaics (listed separately in the DJA) into a full mosaic.
The uncertainty on the stacked flux density is quantified through a bootstrapping technique: being stack the number of input targets in a given band, in each random realisation we re-shuffle the input sample, preserving the same stack by allowing source duplication. We repeat this 100 times, and take the median of the resulting flux distribution as our formal stacked flux. The 1 dispersion around this value is interpreted as the flux error. The rms of the stacked image is computed by stacking the inverse-variance weighted noise map of each target. Our median stacks yield a detection in each band up to 21 m with signal-to-noise444We consider as “noise” the nominal rms of the stacked image; though using the flux error derived via bootstrapping does not change our conclusions. S/N3 (see Table 2).
For ALMA, we perform a stacking analysis in the uv-domain using all publicly available Band-6 observations covering our LRDs in the COSMOS and GOODS-S fields. This analysis is carried out with CASA, following the method described in Wang et al. (2022) (see also Wang et al. 2024; Magnelli et al. 2024), to which we refer for further details of the procedure.
Fig. 4 displays the redshift distribution of our final LRD sample (302 sources, grey dashed), split by field (coloured open histograms). Filled histograms highlight the spectroscopic subsets. Corresponding numbers are listed in Table 1.
@@ Appendix A Stacked samples and fluxes
| PRIMER/COSMOS | NIRCam | [f090w, f115w, f150w, f200w, f277w, f356w, f444w] | 138.9 | 79–80 (7) |
| MIRI | [f770w, f1800w] | 73.0 | 41–42 (3) | |
@@
| MIRI | [f770w, f1800w] | 130.0 | 61–62 (23) |

### v2

@@
Little Red Dots (LRDs) are a puzzling population of extragalactic sources whose origin is highly debated. In this work, we perform a comprehensive stacking analysis of NIRCam, MIRI and ALMA images of a large and homogeneously-selected sample of LRDs from multiple JWST Legacy fields. We report clear evidence for hot-dust emission in the median stacked spectral energy distribution (SED), featuring a rising near-infrared continuum up to rest-frame 3m, which is best explained by a standard dusty AGN structure. Although LRDs are likely a heterogeneous population, our findings suggest that most (50%) LRDs show AGN-heated dust emission, regardless of whether the Optical/Ultraviolet (UV) continuum is stellar or AGN-dominated. In either case, the best-fit dusty-AGN SED, combined with the lack of X-ray detection in the deep Chandra stacks, suggests that Compton-thick (31024 cm-2) gas obscuration is common, and likely confined within the dust sublimation radius (sub0.1 pc). Therefore, we argue that AGN-heated dust does not directly obscure either the Optical/UV continuum or the broad-line region emission, in order to explain the observed blue UV slopes and prominent Balmer features. While a gas-dust displacement is in line with several models, the formation scenario (in-situ or ex-situ) of this pre-enriched hot dust remains unclear.
@@
The James Webb Space Telescope (JWST) has ushered extragalactic astronomy in a transformative era, offering unprecedented access to the high-redshift (4) universe, with the discovery of previously missed Active Galactic Nuclei (AGN; e.g. Matthee et al. 2024). These AGNs are mostly identified via broad emission lines (FWHM1000–2000 km s-1), corresponding to bolometric luminosities Lbol,AGN1044-46 erg s-1 and harboring supermassive black holes (SMBHs, with mass ). Conversely to their lower- analogs, they do not show significant X-ray or radio emission (Ananna et al. 2024, Mazzolari et al. 2025). A non-negligible fraction of these AGN (10–30%; Hainline et al. 2025) feature unique and puzzling properties, and are dubbed as “Little Red Dots” (LRDs, e.g. Matthee et al. 2024, Greene et al. 2024, Labbe et al. 2025). These LRDs are extremely compact (often unresolved at 100 pc scales) and show red optical colors combined with blue Ultraviolet (UV) slopes, resulting in a characteristic “V-shaped” spectral energy distribution (SED), with a sharp turnover near the Balmer break at rest Å (e.g. Setton et al. 2024, Hainline et al. 2025). The vast majority of them are seen at 48, and about 60% of the spectroscopically-confirmed LRDs display broad emission lines (e.g. Greene et al. 2024; Hviding et al. 2025). Notably, LRDs appear to be up to two orders of magnitude more abundant than luminous quasars at z4 (Kokorev et al. 2024, Akins et al. 2024), making them accessible even in narrow-field JWST surveys.
However, an observational consensus of the physical nature of LRDs is still lacking. On the one hand, if LRDs are compact star-forming galaxies (e.g., Labbé et al. 2023), their stellar masses (108-11 M⊙) would imply extreme stellar densities (Pacucci & Narayan 2024). On the other hand, the red optical colors and broad emission lines hint at an AGN origin (Inayoshi et al. 2020, Volonteri et al. 2021, Ji et al. 2025), from which the (e.g. Hα-based) estimates would imply overmassive BHs (/1:100; see e.g. Pacucci et al. 2023; Juodžbalis et al. 2024b; Maiolino et al. 2024; Juodžbalis et al. 2025) compared to local scaling relations (Reines & Volonteri 2015), even without accounting for dust attenuation (e.g. Greene et al. 2025).
In the AGN scenario, the Optical/UV continuum would be moderately reddened by dust (– mag; Casey et al. 2024; Akins et al. 2024; Wang et al. 2025). Yet, this interpretation is apparently at odds with their extreme X-ray and radio weakness implied by the lack of stacked X-ray (Ananna et al. 2024; Yue et al. 2024; Maiolino et al. 2025) and radio emission (Akins et al. 2024; Gloudemans et al. 2025). Recent studies attribute the X-ray weakness to super-Eddington accretion (Madau & Haardt 2024; Lambrides et al. 2024; Pacucci & Narayan 2024; King 2025; Madau 2025; but see e.g. Sacchi & Bogdán 2025).
To also alleviate the tension between the prominent Optical/UV Balmer features and the apparent X-ray weakness, several models postulate the presence of a turbulent, dust-free, and highly dense gas envelope (e.g. Liu et al. 2025; Cenci & Habouzit 2025; Inayoshi & Maiolino 2025; Naidu et al. 2025; Rusakov et al. 2025; de Graaff et al. 2025). These conditions explaining the strong Balmer breaks and the red continua through pure gas opacity, since the gas temperature is too high for the dust to survive. In addition, the broad Balmer lines and other abrorption features can be explained via electron scattering (e.g. Rusakov et al. 2025; Inayoshi et al. 2025) or resonant Balmer scattering (e.g. Naidu et al. 2025). These models mostly differ on the nature of the central object within the gas envelope, which could be either a direct-collapse BH (e.g. Cenci & Habouzit 2025), or a Super-Eddington SMBH (Liu et al. 2025), or a late-stage “Quasi star” (Begelman & Dexter 2025), or a “BH-Star” (e.g. Naidu et al. 2025), or other metal-free objects like a “super-massive star” (Nandal & Loeb 2025).
Such a dust paucity is further supported by the non-detection from deep ALMA stacking, which poses a stringent upper limit to the cold dust budget, corresponding to Mdust106 M⊙ at 6 (Casey et al. 2025; see also Chen et al. 2025; Setton et al. 2025).
In the rest-frame near and mid-IR, a broad diversity of LRD SEDs has been reported in the recent literature. Stacking of MIRI-undetected LRDs reveals a flat SED beyond rest 1.6 m (e.g. Williams et al. 2024; Carranza-Escudero et al. 2025), in contrast to the rising mid-infrared (MIR) continuum expected in obscured AGN, conveying the idea that MIR emission in LRDs is typically dominated by dust-poor stellar populations and not AGN dust. However, based on individual MIRI detections, other studies found that some LRDs display a flat or rising MIR continuum (Pérez-González et al. 2024; Barro et al. 2024; Leung et al. 2024; de Graaff et al. 2025; Ronayne et al. 2025), which is best fitted via composite stellar and AGN templates, suggesting a heterogeneous origin of the IR emission of LRDs.
In this highly debated framework, we caution that previous MIRI-SED studies of LRDs might have been biased towards bright individual objects (e.g. Labbe et al. 2024; Barro et al. 2024; Chen et al. 2025; de Graaff et al. 2025; Wang et al. 2025), or possibly affected either by individual MIRI non-detections (e.g. Pérez-González et al. 2024) or by stacking peculiar LRD sub-samples (e.g. -dropouts; Williams et al. 2024), whose SED could intrinsically differ from that of “typical” LRDs. To explore this possibility, in this paper we present a comprehensive stacking analysis of a large and homogeneously-selected LRD sample with broad NIRCam, MIRI (and ALMA) coverage. Throughout, we assume a Chabrier (2003) IMF and a flat CDM cosmology with = 70 km/s/Mpc, , and .
@@
| PRIMER/COSMOS | NIRCam | [f090w, f115w, f150w, f200w, f277w, f356w, f444w] | 138.9 | 79 (15) |
| MIRI | [f770w, f1800w] | 73.0 | 41 (7) | |
@@
| MIRI | [f770w, f1800w] | 130.0 | 61 (23) | |
| All fields | NIRCam | – | 531.7 | 302 (96) |
| MIRI | – | 292.0 | 163 (51) |
2 Photometric selection of LRDs
We select LRDs from the compilation of Kocevski et al. (2025), consisting of 341 photometrically-selected LRDs in the redshift range 2–11, and retrieved from several JWST/NIRCam surveys. These LRDs were homogeneously selected to display compact morphologies and “V-shaped” continua from spectral slope fitting between blueward and redward of the Balmer break (rest-frame 3645 Å). These criteria enable a self-consistent search for LRD candidates with red optical and blue UV colors over a wide redshift interval (Hainline et al. 2025), although we acknowledge that it does not guarantee an homogeneous physical selection of LRDs.
We replace photometric redshifts () with NIRSpec-based spectroscopic redshifts () taken from NIRSpec Merged Table (version 4222https://zenodo.org/records/15472354; Valentino et al. 2025) of the public Dawn JWST Archive (DJA; Brammer 2023b; de Graaff et al. 2024; Heintz et al. 2025), whenever of high-quality (grade=3, see archive). We augment this list by adding measurements from the JADES public database333https://jades.herts.ac.uk/search/ (for the GOODS-S). This enhances the spectroscopic fraction from 11% (39/341) to 31% (107/341). The mutual agreement between and estimates is very good, with a median absolute deviation (MAD) 1.6%, and only five catastrophic outliers with /(1+)0.2. Hence, we do expect a negligible impact of photo- uncertainties on the median stacks. Fig. 1 displays the redshift distribution of our final LRD sample (302 sources, grey dashed), split by field (coloured open histograms). Filled histograms highlight the spectroscopic subsets. Corresponding numbers are listed in Table 1.
Finally, we restrict ourselves to the NIRCam fields with (at least partial) MIRI coverage in two or more bands (including 18m and/or 21m; see Table 1), which yield the strongest IR-SED constraints. Our final sample counts 302 LRDs taken from the following surveys: the Cosmic Evolution Early Release Science Survey (CEERS; Finkelstein et al. 2023) in the EGS; the JWST Advanced Deep Extragalactic Survey (JADES; Eisenstein et al. 2023) in the GOODS-S, complemented by MIRI imaging from the Systematic Mid-infrared Instrument Legacy Extragalactic Survey (SMILES; Rieke et al. 2024; Alberts et al. 2024); the Public Release IMaging for Extragalactic Research (PRIMER444https://primer-jwst.github.io/) survey in the COSMOS and UDS fields. Out of 302 LRDs, 163 (54%) are covered by MIRI 7.7m, 126 (42%) also by MIRI-18m, and 96 (32%) have a robust . Moreover, we exploit archival ALMA/B6 (1.1 mm) data from the combined A3COSMOS+A3GOODSS surveys (Liu et al. 2019; Adscheid et al. 2024; Magnelli et al. 2024), which are available for a total of 70 LRDs in the COSMOS and GOODS-S fields. The number of sources split down by field, photometric coverage and spectroscopic redshifts is summarized in Table 1.
3 Stacking method and results
Here we outline the input images used for stacking and the stacking method adopted in this work. We perform median stacking in all available NIRCam bands (7) and MIRI bands (2 in PRIMER, 7 in CEERS and 8 in SMILES), at the optical position of all LRDs. We use a fixed pixel scale of 0.06′′, and we extract fluxes from a circular aperture with fixed radius =0.25′′. Image and noise maps are retrieved from the corresponding mosaics of each field (see below). In this step, we discard narrow or medium-band filters (whenever available) to mitigate potential flux boosting in the case of strong emission lines.
We note that stacking is performed in the observed frame, since stacking in the rest-frame would require us to K-correct fluxes by assuming an intrinsic SED, which is instead what we would like to assess, particularly in the rest-frame near-IR (NIR, 1–3m), where an AGN-like or a galaxy-like SED alters the interpretation. However, a rest-frame stacking analysis was carried out in Casey et al. (2024), delivering consistent results with ours. We discuss this comparison in Sect. 4.2.
3.1 Input image maps
For the GOODS-S, we retrieve NIRCam and MIRI images from the JADES and SMILES surveys publicly available on the MAST555SMILES: https://archive.stsci.edu/hlsp/smiles; JADES: https://archive.stsci.edu/hlsp/jades archive. While SMILES maps are already at 0.06′′/px, JADES maps are re-scaled from their native resolution (0.03′′/px) to 0.06′′/px for consistency with the other fields.
For CEERS and PRIMER-UDS, which are unavailable on MAST, we leverage the DJA (Brammer 2023a; Valentino et al. 2023), where the full reduced mosaics are publicly available in each NIRCam and MIRI band666DJA full reduced mosaics in each band are taken from https://s3.amazonaws.com/grizli-v2/JwstMosaics/v7/index.html, all with a pixel scale of 0.04′′ and in units of [10 nJy]. We do a resampling of all mosaics to 0.06′′/px for consistency with SMILES and PRIMER-COSMOS maps, while preserving astrometry and total flux over a given aperture. For PRIMER-UDS, we also merge the East and West mosaics (listed separately in the DJA) into a full mosaic.
For ALMA, instead, we perform mean stacking in the -plane, using all publicly available Band-6 observations covering our LRDs in the COSMOS and GOODS-S fields. Mean stacking is needed in order to homogeneously combine archival data at different frequency, angular resolution and sensitivity. This analysis is carried out with CASA, following the method described in Wang et al. (2022) (see also Wang et al. 2024; Magnelli et al. 2024), to which we refer for further details of the procedure.
3.2 Stacked fluxes
In Fig. 2 we show the stacked cutouts (3′′3′′) in each band, sorted by increasing wavelength. For comparison, the dashed white circle marks the corresponding PSF FWHM. As MIRI images are strongly oversampled, to aid visualization we smooth them with a Gaussian kernel, adopting a larger radius for larger PSF FWHM (see caption of Fig. 2). The stacked ALMA/B6 image is obtained in the -plane and imaged at 0.1′′/px scale. Except for MIRI/f2550w and ALMA/B6, all other median stacks lead to a detection with signal-to-noise ratio (S/N) greater than 3. We stress that all detections are consistent with being unresolved, as their extent is smaller or comparable with the PSF FWHM of the corresponding band (white circles in Fig. 2).
The uncertainty on the stacked flux density is quantified through a bootstrapping technique: being stack the number of input targets in a given band, in each random realisation we re-shuffle the input sample, preserving the same stack by allowing source duplication. We repeat this 100 times, and take the median of the resulting flux distribution as our formal stacked flux. The 1 dispersion around this value is interpreted as the flux error. The rms of the stacked image is computed by stacking the inverse-variance weighted noise map of each target. Our median stacks yield a detection in each band up to 21 m with signal-to-noise777We consider as “noise” the nominal rms of the stacked image; though using the flux error derived via bootstrapping does not change our conclusions. S/N3. All fluxes, errors, rms of the stacked images are listed in Table 2 along with the effective number of sources stacked in each band.
We refer the reader to Appendix A, where we test the robustness of our stacking results against possible biases. In particular, it is worth noting that (rms-weighted) mean stacking delivers systematically higher fluxes than median stacking, by a factor of 2–3 in all bands. In Appendix A.2, we demonstrate that, at least in MIRI bands, these fluctuations are entirely due to individual MIRI detections (dropping from 76% at f770w to 0% at f2100w), without which the mean and median stacked fluxes are matched one another. Therefore, we reiterate that mean stacking is highly sensitive to a few brighter detections, whereas median stacking is more representative of the bulk LRD population.